Shift report 20030116 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Jan 16 2003 - 00:19:05 EST

  • Next message: Pawel Staszel: "DC data status"
    ------------------------------------------------------------
    Date: Thu Jan 16 00:19:05 2003, Shift: 16:00-24:00
    Supervisor: JIJ, with: 
    Log book#13 pages 66 - 66
    
    Shift Summary:
    ------------------------------------------------------------
    1. 
    No physics beam. Currently some low luminosity (6 bunches) beam, but high background. Even with BB coincidences of 200-300/s, H1 and H2 have backgrounds varying from 5k/s to above 100k/s. It seems to stabilize now on 5k/s and 10k/s in H1 and H2, respectively. (MCR just reported they had done some work on the yellow beam.)
    
    2.
    7404 - Pedestal run
    7405 - TPC pedestal run
    (Rings were filled during this run, and therefore not used as input to DAQ.)
    7406 - DVM run
    
    3.
    None! (I repeat none!!!)
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Jan 16 2003 - 00:19:56 EST